### v1

@@ 1 Introduction
This situation has triggered the interest of regulatory agencies and the machine learning community in developing data best practices, such as dataset documentation. Recent public regulatory initiatives (such as the European AI Act [^1] European AI Act required documentation: Annex IV: https://www.euaiact.com/annex/4 and the AI Right of Bills [^2] https://www.whitehouse.gov/ostp/ai-bill-of-rights) and relevant scientific works have provided general guidelines for creating standard dataset documentation (Gebru et al.,, 2021; McMillan-Major et al.,, 2021; Bender and Friedman,, 2018; Holland et al.,, 2020; Micheli et al.,, 2023). More recent works have proposed a structured format for these guidelines (Giner-Miguelez et al., 2023e, ; MLCommons,, 2023), enabling them to be ingested by data search engines like Google Dataset Search (Brickley et al.,, 2019), increasing their discoverability. In these proposals, the authors identify which dimensions, such as the provenance of the dataset or the potential social issues, may influence how the dataset is used and the quality of the ML models trained with it.
@@
To validate our approach, we selected a subset of the papers published in two scientific data journals, Nature’s Scientific Data [^3] https://www.nature.com/sdata/ and Elsevier’s Data in Brief444 https://www.sciencedirect.com/journal/data-in-brief, all describing scientific datasets. First, we manually described these papers in the specified dimensions, and then, we generated automatic descriptions of the papers with our method using two different LLMs (GPT3.5 (Ouyang et al.,, 2022) and FLAN-UL2 (Tay et al.,, 2023)). The results were then reviewed by comparing both descriptions and evaluating the accuracy and faithfulness—i.e., whether the generated answer to the input documents was truthful or not (Maynez et al.,, 2020; Creswell and Shanahan,, 2022). Finally, we present the open-source tool (Giner-Miguelez et al., 2023a, ) implementing our method suited to analyze the documentation of scientific datasets. The tool ingests the dataset documentation (e.g., data papers) and is able to extract the demanded dimensions and check its level of completeness. A public demo of the tool can be found online (Giner-Miguelez et al., 2023c, ), and the experiment’s results and data are available in an open-source repository (Giner-Miguelez et al., 2023d, ).
@@ 4.2 A dataset of annotated data papers
From the selected sample of data papers, and to properly evaluate our method, we created a dataset of ground truth labels describing each of the dimensions we presented in Section 2. To create the labels, the authors have manually described each of the datasets using their data papers according to the dimension presented in Section 2. At least two authors have described each dataset, and, in case of conflicts, the participants have agreed on the final version of the description in subsequent meetings. Each row of the resultant dataset is composed of the DOI and title of the data paper, the specific dimension and sub-dimension of Section 2, and the agreed description from the authors as ground truth labels. The resultant dataset has been released (Giner-Miguelez et al., 2023d, ) in two different formats together with the experiment results to enable future research in this area.
@@ 4.3 Experiment setup and validation
Finally, once we got the results, we evaluated them by comparing both descriptions, the manual we did, and the one generated by the models. We evaluated whether the obtained results were correct or not (accuracy), and in the case of being incorrect, whether the results were truthful with the source document (faithfulness (Maynez et al.,, 2020)). In that sense, erroneous results that were not truthful with the source have been annotated as hallucinations. To ensure the quality of the annotations, at least two authors have evaluated each of the datasets, and the inter-annotation agreement is provided together with the results. A replication package (Giner-Miguelez et al., 2023d, ) has been released along with the annotated dataset, including the source documents and code used in the experiment.
@@ 5.2 Automating the dataset’s discoverability
Furthermore, novel initiatives to improve dataset discoverability and reuse are beginning to emerge in the ML community. Initiatives such as DescribeML (Giner-Miguelez et al., 2023e, ), a domain-specific language to describe datasets, or Croissant (MLCommons,, 2023), a high-level format for describing machine learning datasets based on Schema.org [^11] https://schema.org/, propose to create machine-readable dataset documentation that can be easily indexed by search engines such as the popular Google Dataset Search [^12] https://datasetsearch.research.google.com/. As some of the dimensions these proposals demand are already covered by of our work, our approach could easily be adapted to automatically generate these structured metadata, facilitating the discoverability of well-documented datasets.
@@ 6 Tool support
To facilitate the adoption of our method, we have developed DataDoc Analyzer (Giner-Miguelez et al., 2023a, ), an open-source tool to analyze the documentation of scientific datasets. The tool implements an ML pipeline, including the proposed approach at its core, to extract the demanded dimensions from the dataset documentation and provide a completeness report. The tool is presented with two user interfaces: a demo-based one, intended for test purposes; and an API, ready to be integrated into any data pipeline.
@@
The tool is presented with two distinct user interfaces: a Web UI designed to demonstrate the tool’s capabilities and suitable for analyzing a single document, and an API ready to be integrated into any data processing pipeline. The API comprises endpoints that analyze the documentation for each dimension and return the results in JSON format, ready to be ingested into any data processing pipeline. We implemented the Web UI with Graddio [^13] https://gradio.app/ and the API with FastAPI [^14] https://fastapi.tiangolo.com/. The tool comes with documentation and usage instructions that can be found in the repository, and a docker image (Giner-Miguelez et al., 2023b, ) is provided to facilitate its deployment. Regarding response time, processing unseen documents takes between 50 and 60 seconds. For already-seen documents time goes down to between 20 and 25 seconds for each dimension since the data processing phase is cached. The tool can be found in an open repository (Giner-Miguelez et al., 2023a, ) and a live demo can be found as a space in Huggingface (Giner-Miguelez et al., 2023c, ).

### v2

@@ 1 Introduction
This situation has triggered the interest of regulatory agencies and the machine learning community in developing data best practices, such as dataset documentation. Recent public regulatory initiatives (such as the European AI Act [^1] European AI Act required documentation: Annex IV: https://www.euaiact.com/annex/4 and the AI Right of Bills [^2] https://www.whitehouse.gov/ostp/ai-bill-of-rights) and relevant scientific works have provided general guidelines for creating standard dataset documentation (Gebru et al.,, 2021; McMillan-Major et al.,, 2021; Bender and Friedman,, 2018; Holland et al.,, 2020; Micheli et al.,, 2023). More recent works have proposed a structured format for these guidelines (Giner-Miguelez et al., 2023f, ; MLCommons,, 2023), enabling them to be ingested by data search engines like Google Dataset Search (Brickley et al.,, 2019), increasing their discoverability. In these proposals, the authors identify which dimensions, such as the provenance of the dataset or the potential social issues, may influence how the dataset is used and the quality of the ML models trained with it.
@@
To validate our approach, we selected a subset of the papers published in two scientific data journals, Nature’s Scientific Data [^3] https://www.nature.com/sdata/ and Elsevier’s Data in Brief444 https://www.sciencedirect.com/journal/data-in-brief, all describing scientific datasets. First, we manually described these papers in the specified dimensions, and then, we generated automatic descriptions of the papers with our method using two different LLMs (GPT3.5 (Ouyang et al.,, 2022) and FLAN-UL2 (Tay et al.,, 2023)). The results were then reviewed by comparing both descriptions and evaluating the accuracy and faithfulness—i.e., whether the generated answer to the input documents was truthful or not (Maynez et al.,, 2020; Creswell and Shanahan,, 2022). Finally, we present the open-source tool (Giner-Miguelez et al., 2023a, ) implementing our method suited to analyze the documentation of scientific datasets. The tool ingests the dataset documentation (e.g., data papers) and is able to extract the demanded dimensions and check its level of completeness. A public demo of the tool can be found online (Giner-Miguelez et al., 2023d, ), and the experiment’s results and data are available in an open-source repository (Giner-Miguelez et al., 2023e, ).
@@ 4.2 A dataset of annotated data papers
From the selected sample of data papers, and to properly evaluate our method, we created a dataset of ground truth labels describing each of the dimensions we presented in Section 2. To create the labels, the authors have manually described each of the datasets using their data papers according to the dimension presented in Section 2. At least two authors have described each dataset, and, in case of conflicts, the participants have agreed on the final version of the description in subsequent meetings. Each row of the resultant dataset is composed of the DOI and title of the data paper, the specific dimension and sub-dimension of Section 2, and the agreed description from the authors as ground truth labels. The resultant dataset has been released (Giner-Miguelez et al., 2023e, ) in two different formats together with the experiment results to enable future research in this area.
@@ 4.3 Experiment setup and validation
Finally, once we got the results, we evaluated them by comparing both descriptions, the manual we did, and the one generated by the models. We evaluated whether the obtained results were correct or not (accuracy), and in the case of being incorrect, whether the results were truthful with the source document (faithfulness (Maynez et al.,, 2020)). In that sense, erroneous results that were not truthful with the source have been annotated as hallucinations. To ensure the quality of the annotations, at least two authors have evaluated each of the datasets, and the inter-annotation agreement is provided together with the results. A replication package (Giner-Miguelez et al., 2023e, ) has been released along with the annotated dataset, including the source documents and code used in the experiment.
@@ 5.2 Automating the dataset’s discoverability
Furthermore, novel initiatives to improve dataset discoverability and reuse are beginning to emerge in the ML community. Initiatives such as DescribeML (Giner-Miguelez et al., 2023f, ), a domain-specific language to describe datasets, or Croissant (MLCommons,, 2023), a high-level format for describing machine learning datasets based on Schema.org [^11] https://schema.org/, propose to create machine-readable dataset documentation that can be easily indexed by search engines such as the popular Google Dataset Search [^12] https://datasetsearch.research.google.com/. As some of the dimensions these proposals demand are already covered by of our work, our approach could easily be adapted to automatically generate these structured metadata, facilitating the discoverability of well-documented datasets.
@@ 6 Tool support
To facilitate the adoption of our method, we have developed DataDoc Analyzer (Giner-Miguelez et al., 2023b, ), an open-source tool to analyze the documentation of scientific datasets. The tool implements an ML pipeline, including the proposed approach at its core, to extract the demanded dimensions from the dataset documentation and provide a completeness report. The tool is presented with two user interfaces: a demo-based one, intended for test purposes; and an API, ready to be integrated into any data pipeline.
@@
The tool is presented with two distinct user interfaces: a Web UI designed to demonstrate the tool’s capabilities and suitable for analyzing a single document, and an API ready to be integrated into any data processing pipeline. The API comprises endpoints that analyze the documentation for each dimension and return the results in JSON format, ready to be ingested into any data processing pipeline. We implemented the Web UI with Graddio [^13] https://gradio.app/ and the API with FastAPI [^14] https://fastapi.tiangolo.com/. The tool comes with documentation and usage instructions that can be found in the repository, and a docker image (Giner-Miguelez et al., 2023c, ) is provided to facilitate its deployment. Regarding response time, processing unseen documents takes between 50 and 60 seconds. For already-seen documents time goes down to between 20 and 25 seconds for each dimension since the data processing phase is cached. The tool can be found in an open repository (Giner-Miguelez et al., 2023a, ) and a live demo can be found as a space in Huggingface (Giner-Miguelez et al., 2023d, ).